A stone is thrown downward with an initial velocity of 5m/s from a 20-meter high building. Find the time it takes to reach the ground.
1
Expert's answer
2021-08-31T01:42:55-0400

"h= ut + \frac{1}{2}gt^2\\
20= 5t + \frac{1}{2}*9.8*t^2\\
\mathrm{Multiply\:both\:sides\:by\:}10\\
20\cdot \:10=5t\cdot \:10+\frac{1}{2}\cdot \:9.8t^2\cdot \:10\\
50t+49t^2=200\\
\mathrm{Subtract\:}200\mathrm{\:from\:both\:sides}\\
50t+49t^2-200=200-200\\
\mathrm{Simplify}\\
49t^2+50t-200=0\\
t_{1,\:2}=\frac{-50\pm \sqrt{50^2-4\cdot \:49\left(-200\right)}}{2\cdot \:49}\\
t_1=\frac{-50+10\sqrt{417}}{2\cdot \:49},\:t_2=\frac{-50-10\sqrt{417}}{2\cdot \:49}\\
t=\frac{5\left(\sqrt{417}-5\right)}{49},\:t=-\frac{5\left(5+\sqrt{417}\right)}{49}\\
t=1.57352 s"